# A Wireless AI-Generated Content (AIGC) Provisioning Framework Empowered by Semantic Communication

### v2

@@ III-C Training Algorithm
$loss_{2}=\mathbb{E}_{\mathcal{E}(s),{\epsilon}\sim\mathcal{N}(0,I),t}\left[\|{\epsilon}-{\epsilon}_{\theta}\left(z,t|z_{\scriptscriptstyle t\!e\!x\!t}\right)\|_{2}^{2}\right].$ (18)
@@
IV Intelligent Workload Trade-off Scheme Design
@@
IV-A T2I Service Evaluation Metrics
In our SemAIGC framework, image quality and latency are the two crucial performance metrics that affect service quality. The image quality is primarily influenced by semantic similarity and reconstruction likelihood. We can fetch the image quality score by evaluating the loss between the SemAIGC image and the target image [^2] In practice, if there is no target image available, the aesthetics scores predictor can be used to evaluate the image quality in model training and image generation [29]., as defined in (17). Meanwhile, the latency of accessing T2I service by the local receiver $j$ from edge transmitter $i$ consists of transmission delay $L_{1}^{i,j}$, edge computing delay $L_{2}^{i,j}$, and local computing delay $L_{3}^{i,j}$, i.e., $L_{i,j}=L_{1}^{i,j}+L_{2}^{i,j}+L_{3}^{i,j}$. Let $O_{i,j}$ denote the data size of semantic information transmitted by edge transmitter $i$ to local receiver $j$, then the transmission delay is
$L_{1}^{i,j}=\frac{O_{i,j}}{v_{i,j}},$ (19)
@@ IV-A T2I Service Evaluation Metrics
For the computing delay, according to Amdahl’s law, the computing delay consists of the computing time of the serialized part and the computing time of the parallelizable part [30]. In this work, the computing speed of edge transmitter $i$ is majorly dependent on the GPU [^3] The GPUs work within an ideal environment with suitable device temperature, enough graphics memory, and ample processing memory. . Meanwhile, the majority of the computing workload is caused by the reverse diffusion process of the semantic generation module and fine-tuning module, we denote the delay resulting from other modules as a small constant $l$ [31]. Therefore, the edge computing delay is calculated as
$L_{2}^{i,j}$ $=\frac{(1-\tau)W_{i}(T)O_{s}}{\nu_{i}}+\frac{\tau W_{i}(T)O_{s}}{\nu_{i}m_{i}}+l$ (20)
$=\frac{W_{i}(T)O_{s}}{\nu_{i}}\left(1-\tau+\frac{\tau}{m_{i}}\right)+l,$
where $W_{i}(T)$ is the processing density (in GPU/CPU cycles/bits) of the transmitter UNet that corresponds to the total denoising step number $T$ of the semantic generation module, $O_{s}$ is the input data size, and $\tau\in[0,1]$ stands for the parallelizable fraction of AIGC task. Moreover, $\nu_{i}$ is the processing capability (i.e., the GPU/CPU frequency in cycles/s) of each core of the GPU, and $m_{i}$ denotes the GPU core number of edge transmitter $i$. The local computing delay $L_{3}^{i,j}$ is calculated in the same way.
@@
Assuming identical channel quality and resource allocation under the SemAIGC framework and the edge AIGC framework, to make the SemAIGC framework provide lower service latency than the edge AIGC framework, i.e., $L^{i,j}<L_{\scriptstyle edge}$, the compression rate of the AIGC encoder should satisfy
$\frac{Q_{i,j}}{Q_{s}}<\left|1\!-\!\frac{W_{j}(\bar{T})v_{i,j}\left[C_{1}\nu_{i}\!-\!C_{2}\nu_{j}\right]}{\nu_{i}\nu_{j}\!+\!W_{j}(\bar{T})v_{i,j}\left[C_{1}\nu_{i}\!-\!C_{2}\nu_{j}\right]}\right|,$ (21)
Here, $C_{1}$ and $C_{2}$ are constants, and $10<C_{1}<$ 1, $10<C_{2}<$ 1.
@@
$\frac{O_{i,j}}{v_{i,j}}\!+\!\frac{W_{i}(T)O_{s}}{\nu_{i}}(1-\tau+\frac{\tau}{m_{i}})\!+\!\frac{W_{j}(\bar{T})O_{i,j}}{\nu_{j}}(1-\tau+\frac{\tau}{m_{j}})<$ (22)
$\frac{O_{s}}{v_{i,j}}+\frac{W_{i}(T)O_{s}+W_{j}(\bar{T})O_{i,j}}{\nu_{i}}(1-\tau+\frac{\tau}{m_{i}}).$
@@
$O_{s}\cdot\frac{1}{v_{i,j}}.$
Given that $\frac{Q_{i,j}}{Q_{s}}>0$, and $\tau$, $m_{i}$ and $m_{j}$ are constants, we prove Proposition 2. ∎
IV-B Computing Workload Adaptation Problem Formulation
Considering the dynamic resource availability, channel quality, and service requirements, we should adjust the reverse diffusion step numbers of the semantic generation module and fine-tuning module. We formulate the workload adaptation problem as an MDP problem, and the action, state, reward, and objective function of this problem are defined as follows.
Action: In this work, the edge transmitter acts as the agent to decide its encoding workload. We denote $a\in[0,\hat{T}]$ as the action, which is the reverse diffusion steps taken charge by the edge transmitter’s encoder. Here, $\hat{T}$ is the largest reverse diffusion step for image generation. Meanwhile, the number of denoising steps in the fine-tuning module is set to $\hat{T}-a$.
State: The state of system is denoted as $x=\{\bar{W}_{i},\bar{W}_{j},\overline{\mathit{S\!N\!R}}_{i,j},\bar{\nu}_{i},\bar{\nu}_{j},\bar{B}_{i,j},\bar{L}_{j}\}$, where $\bar{W}_{i}$ and $\bar{W}_{j}$ represent the average processing density of edge transmitter $i$ and local receiver $j$, respectively. Moreover, $\bar{\nu}_{i}$ and $\bar{\nu}_{j}$ are the average processing capability of transmitter $i$ and receiver $j$, $\overline{\mathit{S\!N\!R}}_{i,j}$ denotes the average SNR of the link from transmitter $i$ to receiver $j$, $\bar{B}_{i,j}$ is the average bandwidth allocated to this link, and $\bar{L}_{j}$ represents the average delay requirement of receiver $j$. A D3QN model is sufficient to analysis this state space.
Reward and objective function: In order to provide a required image quality with target latency according to the specific requirements of the receiver, the reward function is set as
$R=\left\{\begin{array}[]{lcr}1,&&{L_{j}<\check{L}_{j}},\\ \frac{\hat{L}_{j}-L_{j}}{\hat{L}_{j}-\check{L}_{j}},&&{\check{L}_{j}<L_{j}<\hat{L}_{j}},\\ 0,&&otherwise.\end{array}\right.$ (24)
Here, $[\check{L}_{j},\hat{L}_{j}]$ is the delay requirement of receiver $j$.
@@ IV-C ROOT Scheme
V Simulation Results and Discussions
@@
Figure 6: Convergence performance for four AIGC generation and transmission framework.
@@
V-A Simulation Settings
In the simulation, a workstation with the GPU RTX A6000 acts as the edge transmitter and a laptop with the GPU of RTX1080 Ti acts as the local receiver. The bandwidth of transmitter provided for receiver does not exceed 20 MHz, and the SNR is set to a range of $[-6,15]dB$. Additionally, the latency requirement of the receiver is within the range of $[5,25]s$. The satisfaction score within $[0,1]$ is issued according to the latency for the receiver to fetch appropriate image quality from transmitter. If the generation time cost of one image exceeds the default maximum number of $60s$, the task is failed.
The model training is composed of two major parts, i.e., SemAIGC model training and ROOT scheme training. For the SemAIGC model, the Stable Diffusion v1-5 model checkpoint is used in the pre-trained model [19]. The corresponding semantic extraction module of text is deployed with pre-trained OpenAI CLIP [21]. Meanwhile, the pre-trained VAE encoder and decoder are used as the image semantic extraction module and autoencoder decoder module, respectively [28]. In this work, to adjust the UNet weights of the fine-tuning module, a subset within the laion-aesthetics v2 5+ is used, which contains high-resolution figures, and corresponding caption texts. In training, all the images in this subset are filtered to a size of $512\times 512\times 3$. All the pre-processed images are extracted as $1\times 4\times 64\times 64$ image semantic information by VAE encoder for UNet training. Moreover, the UNet is trained with 1 middle block in $8\times 8$ resolution, 12 encoding blocks, and 12 decoding blocks in 4 resolutions, i.e., $64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$ [19].
For the D3QN setting in the ROOT scheme, both the DQN network and target network have an input layer with 7 neurons and a 10-neural output layer, as well as 2 hidden layers with each of 256 neurons [32]. We employ $ReLU$ as the activation function between the input layer and the two hidden layers. The training minibatch size is set to 64, and the reward discount rate is set to 0.99.
@@ V-B Numerical Results
Then, we evaluate its convergence performance by comparing its average reward with that of four alternative benchmarks, as shown in Fig. 6. It is worth noting that these four benchmarks employ the same reward function as the SemAIGC framework. We observe that SemAIGC excels in convergence speed, reaching convergence within approximately 350 episodes. This superior performance can be attributed to the ROOT scheme of SemAIGC, which intelligently adjusts the workload for the transmitter and receiver based on resource availability and channel quality.
Fig.7 demonstrates the effectiveness of the ROOT scheme for SemAIGC in terms of intelligently adjusting denoising steps, so as computing workload. From Fig.7(a) and Fig.7(b), we observe that the ROOT scheme of SemAIGC intelligently allocates more workload to the transmitter under batter channel conditions and higher computing resource availability, respectively.
@@
Additionally, Fig. 8 illustrates the average satisfaction score of the user when utilizing the SemAIGC framework, in response to dynamic changes in channel quality and the percentages of available computing resources (PACR) for the transmitter. Notably, the satisfaction score exhibits an increase, ranging from approximately $45\%$ to $98\%$, in tandem with the PACR and SNR. Although under the most challenging conditions, characterized by the poorest channel quality with an SNR of $-6dB$, SemAIGC can still attain an acceptable satisfaction score of approximately $70\%$, provided that the transmitter possesses sufficient computational resources. This impressive performance can be attributed to that the transmitter with ample computing power can speedily generate semantic information, and the fine-tuning module can effectively and promptly mitigate channel noise while preserving valuable semantic information.
@@
To further demonstrate the performance of SemAIGC, we conduct a comparative analysis of its average latency and satisfaction score with other four frameworks, as depicted in Fig. 9. As seen in Fig. 9(a), when computational resources are fixed and the SNR varies dynamically, SemAIGC consistently maintains an average latency below $25s$, surpassing the performance of the other three frameworks. This superiority is even more pronounced in Fig. 9(b). Although SemAIGC exhibits a slightly lower average latency than Edge AIGC by a mere $2s$, its satisfaction rate outperforms Edge AIGC by approximately $40\%$. This is due to the fact that SemAIGC prioritizes delivering images within a required time range rather than blindly pursuing extremely low latency. Furthermore, Fig. 9(c) and Fig. 9(d) offer insights into the performance of these frameworks under varying SNRs. In Fig. 9(c), it becomes evident that under the harsh channel, the user with both the non-fine-tuning AIGC and Edge AIGC framework can experience significantly higher latency, approximately $25s$ and $45s$ greater than that of SemAIGC, respectively. This disparity arises from the absence of UNet, which is responsible for denoising received information from the receiver side and ensuring suitable image quality within the given time constraints. Additionally, from Fig. 9(d), it is more obvious that the other four frameworks cannot provide a satisfying service as the SemAIGC, which further validates the effectiveness and robustness of SemAIGC.
VI Conclusion

### v3

@@ 3.3 Training Algorithm
$loss_{2}=\mathbb{E}_{\mathcal{E}(\mathbf{s}),{\bm{\epsilon}}\sim\mathcal{N}(0,\bm{I}),t}\left[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\bm{\theta}}\left(\mathbf{z},t|\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t}\right)\|_{2}^{2}\right].$ (19)
@@
4 Intelligent Workload Trade-off Scheme Design
@@
4.1 T2I Service Evaluation Metrics
In our SemAIGC framework, image quality and latency are the two crucial performance metrics that affect service quality. The image quality is primarily influenced by semantic similarity and reconstruction likelihood. We can fetch the image quality score by evaluating the loss between the SemAIGC image and the target image [^2] In practical use, if there is no target image available, the aesthetics scores predictors or user feedback mechanisms can be used to evaluate the image quality in model training and image generation [32]., as defined in (18). Meanwhile, the latency of accessing T2I service by the local receiver $j$ from edge transmitter $i$ consists of transmission delay $L_{1}$, edge computing delay $L_{2}$, and local computing delay $L_{3}$, i.e., $L=L_{1}+L_{2}+L_{3}$. Let $O_{i,j}$ denote the data size of semantic information transmitted by edge transmitter $i$ to local receiver $j$, then the transmission delay is
$L_{1}=\frac{O_{i,j}}{v_{i,j}},$ (20)
@@ 4.1 T2I Service Evaluation Metrics
For the computing delay, according to Amdahl’s law, the computing delay consists of the computing time of the serialized part and the computing time of the parallelizable part [33]. In this work, the computing speed of edge transmitter $i$ is majorly dependent on the GPU [^3] It is assumed that the GPUs work within an ideal environment with suitable device temperature, enough graphics memory, and ample processing memory. In practice, real-time monitoring techniques can be integrated into the proposed ROOT for accounting these environmental variations.. Meanwhile, the majority of the computing workload is caused by the reverse diffusion process of the semantic generation module and fine-tuning module, we denote the delay resulting from other modules as a small constant $l$ [34]. Additionally, since the D3QN model within ROOT merely involves a small-scale network, and can pre-make trade-off decisions within a few milliseconds, the additional computational complexity and latency in AIGC services are minimum. Therefore, the edge computing delay is calculated as
$L_{2}$ $=\frac{(1-\tau)W_{i}(T)O_{\mathbf{s}}}{\nu_{i}}+\frac{\tau W_{i}(T)O_{\mathbf{s}}}{\nu_{i}m_{i}}+l$ (21)
$=\frac{W_{i}(T)O_{\mathbf{s}}}{\nu_{i}}\left(1-\tau+\frac{\tau}{m_{i}}\right)+l,$
where $W_{i}(T)$ is the processing density (in GPU/CPU cycles/bits) of the transmitter UNet that corresponds to the total denoising step number $T$ of the semantic generation module, $O_{\mathbf{s}}$ is the input data size, and $\tau\in[0,1]$ stands for the parallelizable fraction of AIGC task. Moreover, $\nu_{i}$ is the processing capability (i.e., the GPU/CPU frequency in cycles/s) of each core of the GPU, and $m_{i}$ denotes the GPU core number of edge transmitter $i$. The local computing delay $L_{3}$ is calculated in the same way.
@@
Assuming identical channel quality and resource allocation under the SemAIGC framework and the edge AIGC framework, to make the SemAIGC framework provide lower service latency than the edge AIGC framework, i.e., $L<L_{\scriptstyle edge}$, the compression rate of the AIGC encoder should satisfy
$\frac{O_{i,j}}{O_{\mathbf{s}}}<\left|1\!-\!\frac{W_{j}(\bar{T})v_{i,j}\left[C_{1}\nu_{i}\!-\!C_{2}\nu_{j}\right]}{\nu_{i}\nu_{j}\!+\!W_{j}(\bar{T})v_{i,j}\left[C_{1}\nu_{i}\!-\!C_{2}\nu_{j}\right]}\right|,$ (22)
where $C_{1}$ and $C_{2}$ are constants, and $10<C_{1}<$ 1, $10<C_{2}<$ 1.
@@
$\frac{O_{i,j}}{v_{i,j}}\!+\!\frac{W_{i}(T)O_{\mathbf{s}}}{\nu_{i}}(1-\tau+\frac{\tau}{m_{i}})\!+\!\frac{W_{j}(\bar{T})O_{i,j}}{\nu_{j}}(1-\tau+\frac{\tau}{m_{j}})<$ (23)
$\frac{O_{\mathbf{s}}}{v_{i,j}}+\frac{W_{i}(T)O_{\mathbf{s}}+W_{j}(\bar{T})O_{i,j}}{\nu_{i}}(1-\tau+\frac{\tau}{m_{i}}).$
@@
$O_{\mathbf{s}}\cdot\frac{1}{v_{i,j}}.$
Given that $\frac{O_{i,j}}{O_{\mathbf{s}}}>0$, and $\tau$, $m_{i}$ and $m_{j}$ are constants, we prove Proposition 2. ∎
4.2 Computing Workload Adaptation Problem Formulation
Let $\hat{L}_{i}$ be the lowest latency that edge AIGC or local AIGC can achieve for generate image in time $i$, $\Psi$ denotes the quality factor of the generated image, if the quality fail to meet the requirement, $\Psi=0$; otherwise, $\Psi=1$. The optimization problem can be formulated as
$\max_{T_{i},\bar{T}_{i}}\sum_{i=1}^{M_{t}}\Psi_{i}(\hat{L}_{i}-L_{i}),$ (25a)
$s.t.\;$ $T_{i}\in\mathbb{Z}_{\geq 0},\forall i,$ (25b)
$\bar{T}_{i}\in\mathbb{Z}_{\geq 0},\forall i,$ (25c)
$T_{i}+\bar{T}_{i}>0,\forall i,$ (25d)
$\Psi\in\{0,1\},\forall i,$ (25e)
where constraints (25b)-(25d) ensure that the denoising steps of the transmitter and receiver are positive integer or zero, and the total denoising step larger than 0. Constraint 25e restricts the quality factor of the generated image.
To solve this problem, we transform it into an MDP problem, and the action, state, reward, and objective function of this problem are defined as follows. Action: In this work, the edge transmitter acts as the agent to decide its encoding workload. We denote $a\in[0,\hat{T}]$ as the action, which is the reverse diffusion steps taken charge by the edge transmitter’s encoder. Here, $\hat{T}$ is the largest reverse diffusion step for image generation. Meanwhile, the number of denoising steps in the fine-tuning module is set to $\hat{T}-a$.
State: The state of system is denoted as $x=\{\bar{W}_{i},\bar{W}_{j},\overline{\mathit{S\!N\!R}}_{i,j},\bar{\nu}_{i},\bar{\nu}_{j},\bar{B}_{i,j},\bar{L}\}$, where $\bar{W}_{i}$ and $\bar{W}_{j}$ represent the average processing density of edge transmitter $i$ and local receiver $j$, respectively. Moreover, $\bar{\nu}_{i}$ and $\bar{\nu}_{j}$ are the average processing capability of transmitter $i$ and receiver $j$, $\overline{\mathit{S\!N\!R}}_{i,j}$ denotes the average SNR of the link from transmitter $i$ to receiver $j$, $\bar{B}_{i,j}$ is the average bandwidth allocated to this link, and $\bar{L}$ represents the average delay requirement of receiver $j$. A D3QN model is sufficient to analysis this state space.
Reward: In order to provide a required image quality with target latency according to the specific requirements of the receiver, the reward function is set as
$R=\left\{\begin{array}[]{lcr}1,&&{L\leq\check{L}},\\ \frac{\eta_{L}\check{L}-L}{\eta_{L}\check{L}_{j}-\check{L}},&&{\check{L}\leq L<\eta_{L}\check{L}},\\ 0,&&otherwise.\end{array}\right.$ (26)
Here, $\check{L}_{j}$ is the delay requirement of receiver $j$, $\eta_{L}>1$ is the latency tolerance factor.
@@ 4.3 ROOT Scheme
5 Simulation Results and Discussions
@@
5.1 Simulation Settings
In the simulation, a workstation with the GPU RTX A6000 acts as the edge transmitter and a laptop with the GPU of GTX1080 acts as the local receiver. The bandwidth budget of the link from the transmitter to the receiver is $20MHz$, and the SNR is set to a range of $[-6,15]dB$. The percentage of available computing resources (PACR) is adjusted by controlling the GPU processing time slot. Additionally, the required latency of the AIGC service is randomly generated within the range of $[5,25]s$, while the maximum latency is twice the minimum latency. The satisfaction score is the average reward in (26) for generating multiple images, in the range $[0,1]$. If the generation time per image exceeds the default value of $60s$, the task is failed.
The model training is composed of two major parts, i.e., SemAIGC model training and ROOT scheme training. For the SemAIGC model, the Stable Diffusion v1-5 model checkpoint is used in the pre-trained model [24]. The corresponding semantic extraction module of text is deployed with pre-trained OpenAI CLIP [25]. Meanwhile, the pre-trained VAE encoder and decoder are used as the image semantic extraction module and autoencoder decoder module, respectively [31]. Moreover, for evaluating the quality of generated image quality, the laion-AI/aesthetic-predictor is used to evaluate the aesthetic score. To adjust the UNet weights of the fine-tuning module [36], a subset within the laion-aesthetics v2 5+ is used, which contains high-resolution images, and corresponding caption texts. In training, all the training images are filtered to a size of $512\times 512\times 3$. All the pre-processed images are extracted as $1\times 4\times 64\times 64$ image semantic information by VAE encoder for UNet training. Moreover, the UNet is trained with 1 middle block in $8\times 8$ resolution, 12 encoding blocks, and 12 decoding blocks in 4 resolutions, i.e., $64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$ [24].
For the D3QN setting in the ROOT scheme, both the DQN network and target network have an input layer with 7 neurons and a 20-neural output layer, as well as 2 hidden layers with each of 256 neurons [35]. We employ $ReLU$ as the activation function between the input layer and the two hidden layers. The training minibatch size is set to 100, and the learning rate is set to 0.0003. For controlling action exploration, the exploration probability is initially set to $1$, and decays $1*10^{-4}$ each episode, while the minimum is $0.01$.
Figure 6: Convergence performance for four AIGC generation and transmission frameworks.
@@ 5.1 Simulation Settings
5.2 Numerical Results
@@
Then, we evaluate its convergence performance by comparing its average reward with that of four alternative benchmarks, as shown in Fig. 6. It is worth noting that these four benchmarks employ the same reward function as the SemAIGC framework. From Fig. 6, we observe that although SemAIGC experiences moderate oscillations due to some randomly generated harsh service requirements, it still converges faster than other schemes, reaching convergence within approximately 350 episodes. This superior performance can be attributed to the ROOT scheme of SemAIGC, which intelligently adjusts the workload for the transmitter and receiver based on resource availability and channel quality. Fig. 7 demonstrates the effectiveness of the ROOT scheme for SemAIGC in terms of intelligently adjusting denoising steps, so as computing workload. From Fig. 7(a) and Fig. 7(b), we observe that the ROOT scheme of SemAIGC intelligently allocates more workload to the transmitter under batter channel conditions and higher computing resource availability, respectively.
@@
Additionally, Fig. 8 illustrates the average satisfaction score of the user when utilizing the SemAIGC framework, in response to dynamic changes in channel quality and the PACR for the transmitter. Notably, the satisfaction score exhibits an increase, ranging from approximately $45\%$ to $98\%$, in tandem with the PACR and SNR. Although under the most challenging conditions, characterized by the poorest channel quality with an SNR of $-6dB$, SemAIGC can still attain an acceptable satisfaction score of approximately $70\%$, provided that the transmitter possesses sufficient computational resources. This impressive performance can be attributed to that the transmitter with ample computing power can speedily generate semantic information, and the fine-tuning module can effectively and promptly mitigate channel noise while preserving valuable semantic information.
@@
To further demonstrate the performance of SemAIGC, we conduct a comparative analysis of its average latency and satisfaction score with other four frameworks, as depicted in Fig. 9. As seen in Fig. 9(a), when computational resources are fixed and the SNR varies dynamically, SemAIGC consistently maintains an average latency below $25s$, surpassing the performance of the other three frameworks. This superiority is even more pronounced in Fig. 9(b). Although SemAIGC exhibits a slightly lower average latency than Edge AIGC by a mere $2s$, its satisfaction score outperforms Edge AIGC by approximately $40\%$. This is due to the fact that SemAIGC prioritizes delivering images within a required time range rather than blindly pursuing extremely low latency. Furthermore, Fig. 9(c) and Fig. 9(d) offer insights into the performance of these frameworks under varying SNRs. In Fig. 9(c), it becomes evident that under the harsh channel, the user with both the non-fine-tuning AIGC and Edge AIGC framework can experience significantly higher latency, approximately $25s$ and $45s$ greater than that of SemAIGC, respectively. This disparity arises from the absence of UNet, which is responsible for denoising received information from the receiver side and ensuring suitable image quality within the given time constraints. Additionally, from Fig. 9(d), it is more obvious that the other four frameworks cannot provide a satisfying service as the SemAIGC, which further validates the effectiveness and robustness of SemAIGC.
6 Conclusion